what is the area of a trapezoid if the top measures 10 feet, the bottom measures 14 foot, and the height is 8 ft.
Answer by checkley77(12844)

[(b1+b2)/2]h=area.
[(14+10)/2]8=24/2*8=12*8=96 ft^2 ans.